Shift summary for owl 11/28/99

Jason Hamm (jhamm@physics.arizona.edu)
Sun, 28 Nov 1999 06:01:40 -0700 (MST)


Shift Summary for owl shift, 11/28/99

On Shift: R. Pordes, J. Hamm

Hours of beam available at all intensities: 7
Hours of beam available at nominal intensity: 7
Hours of E799 data (special runs not included): 7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: None

========================================================================
Runs ended this shift: None

Tapes brought to FCC this shift: None
========================================================================
Accesses during shift: None
========================================================================
Overview of shift: Smooth 799 data taking
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
4:15 Onsplit requests UPF052 on ktev4.

4:30 Onspool error messages: corrupted record for
UPE014 and UPT016. Experts e-mailed.

4:45 Onsplit requests UPM041 on ktev1.

5:00 Delete UPF051 cache to free up space on /ktev
1,2,4e.

7:00 Run 15075 still going at end of shift.

========================================================================
Follow up Notes on Previous Shift

The previous run ended with a "boy, I'm full" message, despite the fact
the tapes were obviously not full. This might be due to the /ktev4g and
/ktev4e disk directories filling up. Copies of the murmur log and syslog
are posted on page 10 of the logbook, for those who are interested.

========================================================================
Items to Follow up on Next Shift

CSI 166: comparator errors, frequency=9,range=2,cap=2

========================================================================